# Towards a Physics Foundation Model

### v1

@@ 2 Related work
In physics, the pursuit of foundation models has largely focused on enhancing the generalization of neural surrogates. Researchers have explored meta-learning (Penwarden et al., 2023) and transfer learning (Subramanian et al., 2023; Goswami et al., 2022) to adapt pretrained models to new PDE systems with fewer data samples. Recently, multi-tasks models trained on multiple physical systems were explored McCabe et al. (2024). However, while both groups demonstrated superior accuracy compared to single-physics models, both opted for finetuning to unseen tasks. These efforts represent important progress, but they still fall short of the “train once, deploy anywhere” paradigm that we envision.
@@ 3.1 Architecture
The neural differentiator (blue dashed box) models the partial derivative ($\frac{\partial X}{\partial t}$) of the physical state with respect to time. $X$ is composed for multiple physical fields (channels), such as pressure, temperature, and velocity. To allow for in-context learning, the differentiator receives multiple time snapshots ($X_{t_{i}-n},...,X_{t_{i}}$) of the physical state. The sample is then tokenized by a single linear transformation across spatial and temporal dimensions, yielding non-overlapping spatiotemporal (tubelet (Arnab et al., 2021)) patches. The size of these patches control the number of spatial and temporal pixels encoded in each token. Absolute positional encodings are added to the patches. The spatiotemporal transformer consists of multiple transformer layers with layer norms and attention across all time and space dimensions, illustrated in Figure 1b. We chose this unified attention mechanism over more computationally efficient factorized approaches to ensure maximum expressivity, allowing the model to capture complex, non-separable phenomena like turbulence and shockwave interactions. Finally, a linear transformation (detokenizer) reverts the spatiotemporal patches into the input space.
@@ 3.2 Datasets
To train a model capable of learning general physical principles, we curated a large and diverse corpus of simulation data, comprising eight distinct datasets listed in Table 1. The combined dataset contains over 2.4 million simulation snapshots, totaling 1.8 TB of data. Our data is sourced from both the publicly available "The Well" benchmark (Ohana et al., 2024) and our own custom simulations, described in detail in Appendix 6.6.
@@ 4.1 Multi-physics learning
Figure 2 presents the average and median mean squared error (MSE) across all test sets. GPhyT demonstrates substantial performance gains in average MSE and achievs a 5× reduction in median MSE compared to UNet and a 29× reduction compared to FNO at equivalent model sizes (size M). Detailed per-dataset losses are provided in SI section 6.2. Remarkably, even our smallest variant (GPhyT) with only 10% of the parameters outperforms the full-sized FNO and UNet when comparing median MSE. This efficiency gain suggests that the spatiotemporal attention mechanism captures cross-scale physical interactions more effectively than traditional convolutional or neural operator approaches. Furthermore, GPhyT exhibits favorable scaling properties with performance gains from the S to XL variants, indicating potential for further improvements with increased model capacity.
@@
4.2 Zero-shot in-context learning
@@ 4.3 Long-range Prediction
Figure 5: (a) Autoregressive rollout prediction for velocity magnitude of GPhyT-XL on the Euler shockwaves with periodic boundary conditions. (b) Average mean squared error (MSE) over 50 time steps for all known physical systems. (c) Average MSE over 50 time steps for the novel unseen physical systems. The results can be best viewed on a high-definition digital monitor.
Figure 5a) illustrates a 50-timestep autoregressive rollout for an Euler shockwave. In the initial phase (t=16), GPhyT-XL maintains high fidelity, accurately capturing the intricate vortex structures and shock fronts of the ground truth. As the rollout progresses, error accumulation becomes visible as a loss of high-frequency detail (t=32). However, the model successfully preserves the global dynamics and physical plausibility of the flow; the large-scale structures and boundary interactions remain stable and coherent. The quantitative evaluation across the known physical systems is shown in Figure 5b). For most physics, GPhyT’s exhibit near-linear error accumulation over 25 timesteps. Exceptions are twophase flow, where changes can occur extremely abrupt, and shear flow, where a chaotic initial system converges to a quasi-static solution.
Extension to out-of-distribution systems (Figure 5c) provides further insight into the model’s generalization mechanisms. Systems with modified boundary conditions maintain error trajectories comparable to their training counterparts. Specifically, Euler with open boundaries tracks the periodic case almost identically, suggesting the model has disentangled bulk dynamics from boundary effects. For entirely novel physics error growth rates are higher, highlighting the need for even more accurate and stable physics foundation models.
@@ Acknowledgments
All training, evaluation and plotting code will be published. Furthermore, model checkpoints will be released on Hugging Face. Finally, the datasets which are not available on The Well will be avaiable to download.

### v2

@@ 2 Related work
In physics, the pursuit of foundation models has largely focused on enhancing the generalization of neural surrogates. Researchers have explored meta-learning (Penwarden et al., 2023) and transfer learning (Subramanian et al., 2023; Goswami et al., 2022) to adapt pretrained models to new PDE systems with fewer data samples. Recently, multi-tasks models trained on multiple physical systems were explored McCabe et al. (2024). However, while these groups demonstrated superior accuracy compared to single-physics models, all opted for finetuning to unseen tasks. These efforts represent important progress, but they still fall short of the “train once, deploy anywhere” paradigm that we envision.
@@ 3.1 Architecture
The neural differentiator (blue dashed box) models the partial derivative ($\frac{\partial X}{\partial t}$) of the physical state with respect to time. $X$ is composed of multiple physical fields (channels), such as pressure, temperature, and velocity. To allow for in-context learning, the differentiator receives multiple time snapshots ($X_{t_{i}-n},...,X_{t_{i}}$) of the physical state. The sample is then tokenized by a single linear transformation across spatial and temporal dimensions, yielding non-overlapping spatiotemporal (tubelet (Arnab et al., 2021)) patches. The size of these patches control the number of spatial and temporal pixels encoded in each token. Absolute positional encodings are added to the patches. The spatiotemporal transformer consists of multiple transformer layers with layer norms and attention across all time and space dimensions, illustrated in Figure 1b. We chose this unified attention mechanism over more computationally efficient factorized approaches to ensure maximum expressivity, allowing the model to capture complex, non-separable phenomena like turbulence and shockwave interactions. Finally, a linear transformation (detokenizer) reverts the spatiotemporal patches into the input space.
@@ 3.2 Datasets
To train a model capable of learning general physical principles, we curated a large and diverse corpus of simulation data, comprising seven distinct datasets listed in Table 1. The combined dataset contains over 2.4 million simulation snapshots, totaling 1.8 TB of data. Our data is sourced from both the publicly available "The Well" benchmark (Ohana et al., 2024) and our own custom simulations, described in detail in Appendix 6.6.
@@ 4.1 Multi-physics learning
Figure 2 presents the average and median mean squared error (MSE) across all test sets. GPhyT demonstrates substantial performance gains in average MSE and achievs a 5× reduction in median MSE compared to UNet and a 29× reduction compared to FNO at equivalent model sizes (size M). Detailed per-dataset losses are provided in SI section 6.2. Remarkably, even our smallest variant (GPhyT-S) with only 10% of the parameters outperforms the full-sized FNO and UNet when comparing median MSE. This efficiency gain suggests that the spatiotemporal attention mechanism captures cross-scale physical interactions more effectively than traditional convolutional or neural operator approaches. Furthermore, GPhyT exhibits favorable scaling properties with performance gains from the S to XL variants, indicating potential for further improvements with increased model capacity.
@@
4.2 In-context learning
@@ 4.3 Long-range Prediction
Figure 5: (a) Autoregressive rollout prediction for velocity magnitude of GPhyT-XL on the Euler shockwaves with periodic boundary conditions. (b) Average mean squared error (MSE) over 50 time steps for all known physical systems. (c) Average MSE over 50 time steps for the novel unseen physical systems.
Figure 5a) illustrates an autoregressive rollout for an Euler shockwave with known periodic boundary conditions, exhibiting low error accumulation. In the initial phase (t=16), GPhyT-XL maintains high fidelity, accurately capturing the intricate vortex structures and shock fronts of the ground truth. As the rollout progresses, error accumulation becomes visible as a loss of high-frequency detail (t=32). However, the model successfully preserves the global dynamics and physical plausibility of the flow; the large-scale structures and boundary interactions remain stable and coherent. Similar trajectories for additional systems are presented in Appendix 6.9. The quantitative evaluation across the known physical systems is shown in Figure 5b). For most physics, GPhyT’s exhibit near-linear error accumulation. Exceptions are twophase flow, where changes can occur extremely abrupt, and shear flow, where a chaotic initial system converges to a quasi-static solution.
Extension to out-of-distribution systems (Figure 5c) provides further insight into the model’s generalization mechanisms. Systems with modified boundary conditions maintain error trajectories comparable to their training counterparts. Specifically, Euler with open boundaries tracks the periodic case almost identically, suggesting the model has disentangled bulk dynamics from boundary effects. Similarly, for entirely novel physics error growth rates are higher, highlighting the need for even more accurate and stable physics foundation models.
@@ Acknowledgments
All training, evaluation and plotting code will be published. Furthermore, model checkpoints will be released on Hugging Face. Finally, the datasets which are not available on The Well will be available to download.